40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4526     Entry time: Thu Apr 14 19:05:17 2011
Author: Koji 
Type: Update 
Category: LSC 
Subject: New (temporary) LSC screen 

[Jenne Koji]

The PD signals are transmitted to the suspension now.

The trigger thresholds were set to -1. This means the triggers are always on.

Attachment 1: temporary_LSC_screen.png  40 kB  | Hide | Hide all
temporary_LSC_screen.png
ELOG V3.1.3-